Shift Summary for Day shift, 12/22/99 (Wed.)

Patricia McBride (mcbride@fnpsph)
Wed, 22 Dec 1999 15:41:02 CST


Shift Summary for Day shift, 12/22/99 (Wed.)

On Shift: Patty McBride, Sasha L

Hours of beam available at all intensities: 7.5
Hours of beam available at nominal intensity: 7.0
Hours of E799 data (special runs not included): 6.75
Hours of special runs (laser scans, muon runs...):

Accounting of lost time & intensity:

some KFRENDs, pipeline resets, work on Trigger crate 8, beam tuning
and a fresh start.

========================================================================
Runs ended this shift:

15267 -- BEAM_799 1.47 M events
15268 -- BEAM_799 long run

Tapes brought to FCC this shift:

None.

========================================================================
Accesses during shift:

None

========================================================================
Overview of shift:

Many small startup problems, but we continued to take data
throughout the shift.

Many "HCC inefficient" warnings caused by a high HCC threshold
on CsI channel 1767.

========================================================================
Description of Shift:

15267 07:00 Run ends with ``L3 accept fraction'' warnings and
KTeV2 drops out of the readout. Dart_fresh_start...

15268 07:33 Start of run 15268

07:55 V' latches are in the wrong slices

08:00 Onsplit requests UPF061; Could not find any UPF tapes
so we were forced to label some tapes first!

09:00 SEM is reading 1.23e13
We called MCR and they said that they were
tuning and things would be unstable. SEM reading
reach more than 3e13 at one point, but livetime
indicates it was actually closer to 7e12.

70 09:10 Add 1ns to VV'RF; latches are still jumping around
(userlatch set to 1)

80 Pipeline problems; subtract 1ns from VV' RF
(userlatch set to 2)

88 Many kfrends are required
+2 ns to VV' RF (userlatch 3); V' is jumping
09:55 -2 ns to VV' RF (userlatch 4) VV' latches are stable
from this moment

10:45 Onsplit requests UPK032

13:10 Onsplit requests UPH044

323 14:50 Trigger crate 8 is put back in the readout
after swapping all modules.

350 15:25 Trigger crate 8 problems return; kfrend again
no errors after that

KTeV Logbook XXIII, pages 112-113.

========================================================================
Follow up Notes on Previous Shift

We now have two working flashlights.

========================================================================
Items to Follow up on Next Shifts

There seems to be a problem with the temperature in the blockhouse. The
top of the CsI is too cold. Ed suspects the heaters are not working.

Page Ed when Run 15268 finishes